# Analog Quantum Simulation of Coupled Electron-Nuclear Dynamics in Molecules

### v1

@@ Abstract
Understanding the coupled electron-nuclear dynamics in molecules induced by light-matter interactions is crucial for potential applications of photochemical processes, but it is challenging due to the high computational costs of exact quantum dynamics simulations. Quantum computing has the potential to reduce the computational cost required for exact quantum dynamics simulations by exploiting the quantum nature of the computational device. However, existing quantum algorithms for coupled electron-nuclear dynamics simulation either require fault-tolerant devices, or use the Born-Oppenheimer (BO) approximation and a truncation of the electronic basis. In this work, we present the first analog quantum simulation approach for molecular vibronic dynamics in a pre-BO framework, i.e. without the separation of electrons and nuclei, by mapping the molecular Hamiltonian to a device with coupled qubits and bosonic modes. We show that our approach has exponential savings in resource and computational costs compared to the equivalent classical algorithms. The computational cost of our approach is also exponentially lower than existing BO-based quantum algorithms. Furthermore, our approach has a much smaller resource scaling than the existing pre-BO quantum algorithms for chemical dynamics. The low cost of our approach will enable an exact treatment of electron-nuclear dynamics on near-term quantum devices.
@@ I Introduction
Light-matter interactions are the source of many phenomena in molecular systems such as vision [1, 2, 3], photosynthesis [4, 5, 6, 7], photovoltaics [8, 9, 10, 11], and photocatalysis [12, 13]. In order to apply light-matter interactions to the development of high-functional molecular devices, it is important to understand their underlying mechanisms. In principle, the molecular time-dependent Schrödinger equation can be exactly solved for a system of electrons and nuclei to simulate molecular dynamics. However, it is practically impossible for molecules with more than a few atoms due to the exponential scaling of its computational cost with respect to system size.
@@
The Born-Oppenheimer (BO) approximation neglects the coupling between electronic and nuclear degrees of freedom, known as nonadiabatic coupling (NAC) [14]. Molecular dynamics within in the BO approximation are reduced to nuclear dynamics on a single BO state, i.e. on an eigenstate of electronic Hamiltonian. Although the BO approximation has been successfully applied to study ground-state chemical reactions, it fails to describe light-induced chemical reactions where the molecule can reach strong NAC regions after photo-excitation of the molecular electronic state [14]. The BO approximation can be extended to the group BO approximation (GBOA), which takes into account NACs between a group of BO states considered to be relevant to the dynamics, while NACs to states outside of the group are neglected [14]. The molecular dynamics are then described as nuclear dynamics on multiple BO potential energy surfaces coupled via NACs. The majority of existing simulation methods make use of the GBOA since it extends the scope of molecular dynamics simulation while using conventional electronic structure methods [15, 16, 17]. However, the GBOA can lead to the wrong description of molecular dynamics due to the truncation of the electronic basis set to a small number of BO states. Furthermore, since existing simulation methods with the GBOA require an accurate calculation of BO states and their gradients (including NACs), the computational cost for accurate simulations can become intractable.
@@
Pre-Born-Oppenheimer (pre-BO) methods, on the other hand, naturally include nonadiabatic effects since they treat nuclei and electrons without separation [18, 19, 20, 21, 22, 23, 24]. There have been many studies on the pre-BO theory with both first quantization [18, 19, 20, 21] and second quantization [22, 23, 24] of the molecular Hamiltonian. One well-known approach using the pre-BO framework is the nuclear-electronic orbital (NEO) method which treats certain nuclei as quantum particles using a basis of nuclear orbitals [22]. Recently, an approach based on the multi-configurational time-dependent Hartree method using a second quantization representation of electronic degrees of freedom was developed, which enables the description of coupled electron-nuclear dynamics without electronic potential energy surfaces [24]. Nevertheless, the development of methods in the pre-BO framework is still far behind the methods in the BO framework since the computational cost of a pre-BO treatment of molecular dynamics is much greater.
@@
Quantum computing can significantly reduce the computational cost of the simulation of quantum mechanical systems by exploiting the intrinsic quantum nature of the computational device [25, 26, 27, 28]. Most near-term quantum computing research in the field of quantum chemistry is based on the variational quantum eigensolver (VQE) method [29, 30, 31, 32, 33]. VQE and similar methods typically use the BO approximation to obtain electronic properties at fixed classical nuclear configurations. Although there are several quantum algorithms for chemical dynamics without the BO approximation [34, 35, 36], most methods still adopt the GBOA for near-term applications with variational quantum algorithms [37, 38] or analog mappings [39, 40, 41]. There have been proposals for near-term pre-BO eigenvalue optimization on quantum computers using NEO and VQE [42], as well as proposals for time-dependent algorithms with a real-space, first quantized wavefunction ansatz for electrons and nuclei on fault-tolerant quantum computers [34, 36]. The intractability of quantum dynamics simulations with a pre-BO wavefunction on classical computers suggests that a pre-BO quantum simulation could show an earlier quantum advantage than methods using the BO framework. Therefore, it is desirable to develop near-term, pre-BO quantum algorithms with scaling advantages in quantum resource and operations over existing methods for chemical dynamics.
@@
In this work, we propose the first analog quantum simulation method for molecular vibronic dynamics in the pre-BO framework. Our approach maps the second quantized pre-BO representation of the molecular vibronic Hamiltonian onto an analog quantum device, and thus treats the electron-nuclear interactions exactly. Specifically, we map the nuclear vibrational motions to bosonic modes of a device and use a fermion-qubit mapping for electronic degrees of freedom. We show that our method can efficiently simulate the exact molecular vibronic dynamics within a given single-particle basis set, and therefore it is suitable for accurate near-term quantum simulations of coupled electron-nuclear dynamics. As an example, we show how a single-mode vibronic model system could be simulated on an trapped ion quantum computer on the scale of current devices. Finally, we show how the scaling of our approach out-performs all previous approaches in terms of hardware resources, implementation cost, and classical pre-calculation.
II Theory
@@
Our simulation approach is restricted to vibronic (vibrational + electronic) internal degrees of freedom of the molecule, meaning the translational and rotational degrees of freedom of the atomic nuclear coordinates are removed. In general, this can be achieved by any unitary transformation of the Cartesian atomic coordinates that separates the 6 collective translations and rotations of the molecule, given by vectors $\mathbf{Q}_{\mathrm{trans}}$ and $\mathbf{Q}_{\mathrm{rot}}$, from the $63N_{\mathrm{at}}-$ 6 vibrational internal coordinates, $\mathbf{Q}\equiv\mathbf{Q}_{\mathrm{vib}}$. For the remainder of the manuscript, we assume that mass-weighted normal mode coordinates are used, whereby the translations and rotations are easily identified as the zero-frequency modes. The electronic coordinates are defined as positions of electrons relative to the nuclear center of mass with a fixed orientation, i.e. in the Eckart frame [43]. The Coriolis coupling (between nuclear vibrations and rotations) is excluded. While the removal of molecular rotations is an approximation [43], it is an appropriate one for ultrafast (fs–ps) chemical dynamics simulations, and we provide further justification in the Discussion.
@@
where $\mathbf{r}_{i}$ is the position of the electron $i$ in the Eckart frame, $Q_{\nu}$ is the normal mode coordinate for mode $\nu$, $P_{\nu}=-i\partial/\partial Q_{\nu}$ is the corresponding momentum, $N_{\mathrm{mode}}$ is the number of vibrational modes, $N_{\mathrm{e}}$ is the number of electrons, and $v_{\mathrm{en}}$, $v_{\mathrm{ee}}$, and $V_{\mathrm{nn}}$ are electron-nuclear, electron-electron, and nuclear-nuclear interaction potentials, respectively. We use atomic units ($\hbar=m_{\mathrm{e}}=e=4\pi\epsilon_{0}=1$) here and throughout this paper.
@@
We express the molecular vibronic wavefunction using Slater determinants of the position-dependent spin orbitals for electronic degrees of freedom and Hartree products of harmonic oscillator eigenstates of normal modes for vibrational degrees of freedom. In the second quantization representation, the Slater determinants and the Hartree product can be expressed as occupation number vectors (ONVs). An electronic ONV for a Slater determinant is written as $|\mathbf{n}\rangle_{\mathrm{e}}=|n_{1}\cdots n_{N_{\mathrm{o}}}\rangle_{\mathrm{e}}$, where $n_{j}$ is the occupation of the $j$-th spin orbital and $N_{\mathrm{o}}$ is the number of spin orbitals, while a vibrational ONV for a Hartree product is given as $|\mathbf{v}\rangle_{\mathrm{n}}=|v_{1}\cdots v_{N_{\mathrm{mode}}}\rangle_{\mathrm{n}}$, where $v_{\nu}$ is the occupation (Fock state) of mode $\nu$. Therefore, our wavefunction ansatz becomes
@@
Sasmal and Vendrell [24] showed that above ansatz transforms the molecular vibronic Hamiltonian to the second quantization representation [24]:
@@
where $h_{pq}$, $d_{\nu,pq}$, $g_{\nu,pq}$, and $v_{pqrs}$ are electron integrals whose definitions are given in the Supplemental Material. The $\mathbf{Q}$-dependence of spin-orbitals and Slater determinants are absorbed into the electronic integrals, and the couplings between nuclear and electronic degrees of freedom appear as products of the electron integrals and fermionic ladder operators in the Hamiltonian. The terms akin to nonadiabatic coupling terms in electronic structure ($d_{\nu,pq}$ and $g_{\nu,pq}$) are introduced because the nuclear kinetic energy operator acts on the spin orbital basis functions [24]. The mass-weighted normal mode position $\hat{Q}_{\nu}=\sqrt{1/2\omega_{\nu}}(\hat{b}^{\dagger}_{\nu}+\hat{b}_{\nu})$ and the conjugate momentum $\hat{P}_{\nu}=i\sqrt{\omega_{\nu}/2}(\hat{b}^{\dagger}_{\nu}-\hat{b}_{\nu})$ operators for mode $\nu$ can be expressed in terms of the bosonic ladder operators $\{\hat{b}_{\nu}\}$ and $\{\hat{b}^{\dagger}_{\nu}\}$ where $\omega_{\nu}$ is the (approximate) harmonic frequency of the normal mode $\nu$. Therefore, the terms depending on the normal mode coordinates in Eq. 3 can be expressed in terms of the bosonic ladder operators by using a Taylor series expansion about a reference geometry $\mathbf{Q}_{0}$ up to a reasonable order.
@@
For fixed Taylor expansion orders, $\omega_{\nu}$ and $\mathbf{Q}_{0}$ only affect the accuracy of our approach by the number of harmonic eigenstates required to describe the wavefunction. In other words, a quantum simulator with lower oscillator noise can tolerate lower-accuracy parameters, including approximations derived from the electronic integral Taylor expansions. For practical cases in the near term, we expect that density functional theory (DFT) will provide an ideal balance between accuracy and classical computational cost when finding $\mathbf{Q}_{0}$ and $\omega_{\nu}$, and Kohn-Sham orbitals provide a compact spin orbital basis [44].
@@
In general, the orbital vibronic coupling functions $d_{\nu,pq}$ and $g_{\nu,pq}$ can be extremely localized and non-analytical at some nuclear geometries. This requires a high-order expansion in terms of the bosonic ladder operators in Eq. 3, which greatly increases the number of terms in the Hamiltonian. One way to reduce the expansion order would be a unitary transformation of orbitals for which the derivative couplings vanish, which must satisfy $\partial\mathbf{c}/\partial Q_{\nu}+\mathbf{d}_{\nu}\mathbf{c}=0$, where $\mathbf{c}$ is the transformation matrix to the resulting set of “diabatic” orbitals, $\eta_{m}=\sum_{l}c_{lm}\psi_{l}$. This transformation is identical in form to the many-body BO state diabatization condition [14]. A set of strictly diabatic states does not exist in general for multi-electron states in multi-mode systems [45], and the same arguments apply for orbitals. However, because our goal is simply to achieve a low-order expansion in terms of $\mathbf{Q}$, we can use “quasi-diabatic” orbitals where the above strict diabatization condition is not satisfied. These orbitals can likewise borrow techniques from BO state diabatization [46].
@@
Figure 1: Illustration of analog quantum simulation approaches for molecular vibronic dynamics using trapped-ion architectures. a, The electronic wavefunction is given in terms of spin orbitals. Each diabatic spatial orbital, which depends on the nuclear positions, yields two spin orbitals with opposite spins: $|\phi_{1}\rangle$ and $|\phi_{2}\rangle$ (green), and $|\phi_{3}\rangle$ and $|\phi_{4}\rangle$ (yellow). Only the motion of a single ion (grey) is shown. b, Molecular vibronic dynamics are simulated based on the pre-BO representation with interactions among electronic (cyan) and nuclear (grey) degrees of freedom. c, In this work, we propose a coupled multi-qubit boson (cMQB) approach which maps occupation numbers $\{n_{p}\}$ of spin orbitals $\{|\phi_{p}\rangle\}$ to qubits and nuclear motions to bosonic modes coupled to the qubits, and thus can perform pre-BO dynamics quantum simulations with a complete electronic basis within a given orbital basis set. d, Existing quantum/classical algorithms rely on BO states $\{|\Phi_{i}\rangle\}$ with potential energy surfaces $\{|E_{i}\rangle\}$ found by solving the electronic time-independent Schrödinger equation using the orbital basis at fixed nuclear positions prior to the simulation. e, Molecular vibronic dynamics are solved in the BO framework by propagating the nuclear wavefunction (red and blue densities) on a truncated set of coupled BO states. The truncation leads to a deviation from exact results shown in b.
II.1 Analog quantum simulation

### v2

@@ Abstract
Quantum computing has the potential to reduce the computational cost required for quantum dynamics simulations. However, existing quantum algorithms for coupled electron-nuclear dynamics simulation either require fault-tolerant devices, or use the Born-Oppenheimer (BO) approximation and a truncation of the electronic basis. We present the first analog quantum simulation approach for molecular vibronic dynamics in a pre-BO framework, i.e. without the separation of electrons and nuclei, by mapping the molecular Hamiltonian to a device with coupled qubits and bosonic modes. We show that our approach has exponential savings in resource and computational costs compared to the equivalent classical algorithms. The computational cost of our approach is also exponentially lower than existing BO-based quantum algorithms. Furthermore, our approach has a much smaller resource scaling than the existing pre-BO quantum algorithms for chemical dynamics. The low cost of our approach will enable an exact treatment of electron-nuclear dynamics on near-term quantum devices.
@@ I Introduction
Light-matter interactions are the source of many phenomena in molecular systems such as vision [1, 2, 3], photosynthesis [4, 5, 6, 7], photovoltaics [8, 9, 10, 11], and photocatalysis [12, 13]. In order to apply light-matter interactions to the development of high-functional molecular devices, it is important to understand their underlying mechanisms. In principle, the molecular time-dependent Schrödinger equation can be exactly solved for systems of electrons and nuclei to simulate molecular dynamics. However, such a solution is practically impossible for molecules with more than a few atoms due to the exponential scaling of the computational cost with respect to system size.
@@
The Born-Oppenheimer (BO) approximation neglects the coupling between electronic and nuclear degrees of freedom, known as nonadiabatic coupling (NAC) [14]. While it is useful for properties of the ground electronic state, it fails to describe light-induced chemical reactions where the molecule can reach strong NAC regions after photo-excitation of the molecular electronic state [14]. The BO approximation can be extended to the group BO approximation (GBOA), which takes into account NACs between a group of BO states considered to be relevant to the dynamics, while NACs to states outside of the group are neglected [14]. The molecular dynamics are then described as nuclear dynamics on multiple BO potential energy surfaces coupled via NACs. The majority of existing simulation methods make use of the GBOA since it extends the scope of molecular dynamics simulation while using conventional electronic structure methods [15, 16, 17]. However, truncation of the electronic basis by the GBOA requires some degree of intuition, and it necessarily neglects couplings to higher-lying BO states that alter the molecular dynamics. Furthermore, since existing simulation methods with the GBOA require an accurate calculation of BO states and their gradients (including NACs), the computational cost for accurate simulations can become intractable.
@@
Pre-Born-Oppenheimer (pre-BO) methods, on the other hand, naturally include nonadiabatic effects since they treat nuclei and electrons without separation [18, 19, 20, 21, 22, 23, 24]. There have been many studies on pre-BO theory with both first quantization [18, 19, 20, 21] and second quantization [22, 23, 24] of the molecular Hamiltonian. One approach using the pre-BO framework is the nuclear-electronic orbital (NEO) method, which expands certain nuclei in a basis of nuclear orbitals and those the nuclei as quantum particles [22]. Recently, an approach was developed based on the multi-configurational time-dependent Hartree method using a second quantization representation of electronic degrees of freedom, which enables the description of coupled electron-nuclear dynamics without electronic potential energy surfaces [24]. Nevertheless, the development of methods in the pre-BO framework is still far behind the methods in the BO framework since the computational cost of a pre-BO treatment of molecular dynamics is much greater.
@@
Quantum computing can significantly reduce the computational cost of the simulation of quantum mechanical systems by exploiting the intrinsic quantum nature of the computational device [25, 26, 27, 28]. Most near-term quantum computing research in the field of quantum chemistry has focused on obtaining electronic properties at fixed nuclear configurations based on the variational quantum eigensolver (VQE) method [29, 30, 31, 32, 33] using the BO approximation, with proposed extensions to pre-BO eigenvalues using NEO [34]. For chemical dynamics, although there are several proposed quantum algorithms without the BO approximation for fault-tolerant devices [35, 36, 37], most methods still adopt the GBOA for near-term applications with variational quantum algorithms [38, 39, 40] or analog mappings [41, 42, 43, 44]. The intractability of quantum dynamics simulations with a pre-BO wavefunction on classical computers suggests that a pre-BO quantum simulation could show an earlier quantum advantage than methods using the BO framework.
@@
In this work, we propose the first analog quantum simulation method for molecular vibronic dynamics in the pre-BO framework. Our approach maps the second quantized pre-BO representation of the molecular vibronic Hamiltonian onto an analog quantum device, and thus treats the electron-nuclear interactions exactly. Specifically, we map the nuclear vibrational motions to bosonic modes of a device and use a fermion-qubit mapping for electronic degrees of freedom. We show that our method can efficiently simulate the exact molecular vibronic dynamics within a given single-particle basis set, and therefore it is suitable for accurate near-term quantum simulations of coupled electron-nuclear dynamics. As an example, we show how a single-mode vibronic model system could be simulated on currently available trapped ion quantum computers with existing experimental techniques and noise. Finally, we show how the scaling of our approach out-performs all previous approaches in terms of hardware resources, implementation cost, and classical pre-calculation.
II Results
II.1 Theory
@@
Our simulation approach is restricted to vibronic (vibrational + electronic) internal degrees of freedom of the molecule, meaning the translational and rotational degrees of freedom of the atomic nuclear coordinates are removed. In general, this can be achieved by any unitary transformation of the Cartesian atomic coordinates that separates the 6 collective translations and rotations of the molecule, given by vectors $\mathbf{Q}_{\mathrm{trans}}$ and $\mathbf{Q}_{\mathrm{rot}}$, from the $63N_{\mathrm{at}}-$ 6 vibrational internal coordinates, $\mathbf{Q}\equiv\mathbf{Q}_{\mathrm{vib}}$. For the remainder of the manuscript, we assume that mass-weighted normal mode coordinates are used, whereby the translations and rotations are easily identified as the zero-frequency modes. The electronic coordinates are defined as positions of electrons relative to the nuclear center of mass with a fixed orientation, i.e. in the Eckart frame [45]. The Coriolis coupling (between nuclear vibrations and rotations) is excluded. While the removal of molecular rotations is an approximation [45], it is appropriate for ultrafast (fs–ps) chemical dynamics simulations, as is further justified in the Discussion.
@@
where $\mathbf{r}_{i}$ is the position of the electron $i$ in the Eckart frame, $\nabla_{i}=\partial/\partial\mathbf{r}_{i}$ is its gradient, $Q_{\nu}$ is the normal mode coordinate for mode $\nu$, $P_{\nu}=-i\partial/\partial Q_{\nu}$ is the corresponding momentum, $N_{\mathrm{mode}}$ is the number of vibrational modes, $N_{\mathrm{e}}$ is the number of electrons, and $v_{\mathrm{en}}$, $v_{\mathrm{ee}}$, and $V_{\mathrm{nn}}$ are electron-nuclear, electron-electron, and nuclear-nuclear interaction potentials, respectively. We use atomic units ($\hbar=m_{\mathrm{e}}=e=4\pi\epsilon_{0}=1$) here and throughout this paper.
@@
We express the molecular vibronic wavefunction using Slater determinants of the position-dependent spin orbitals for the electronic degrees of freedom and Hartree products of harmonic oscillator eigenstates of each normal mode for the vibrational degrees of freedom. In the second quantization representation, the Slater determinants and the Hartree product can be expressed as occupation number vectors (ONVs). An electronic ONV for a Slater determinant is written as $|\mathbf{n}\rangle_{\mathrm{e}}=|n_{1}\cdots n_{N_{\mathrm{o}}}\rangle_{\mathrm{e}}$, where $n_{j}$ is the occupation of the $j$-th spin orbital and $N_{\mathrm{o}}$ is the number of spin orbitals, while a vibrational ONV for a Hartree product is given as $|\mathbf{v}\rangle_{\mathrm{n}}=|v_{1}\cdots v_{N_{\mathrm{mode}}}\rangle_{\mathrm{n}}$, where $v_{\nu}$ is the occupation (Fock state) of mode $\nu$. Therefore, our wavefunction ansatz becomes
@@
Sasmal and Vendrell showed that the ansatz above transforms the molecular vibronic Hamiltonian to the second quantization representation [24]:
@@
where $h_{pq}$, $d_{\nu,pq}$, $g_{\nu,pq}$, and $v_{pqrs}$ are electron integrals whose definitions are given in the Supplemental Material [46]. The $\mathbf{Q}$-dependence of spin-orbitals and Slater determinants are absorbed into the electronic integrals, and the couplings between nuclear and electronic degrees of freedom appear as products of the electron integrals and fermionic ladder operators in the Hamiltonian. The orbital vibronic coupling terms (depending on $d_{\nu,pq}$ and $g_{\nu,pq}$) appear because the nuclear kinetic energy operator acts on the spin orbital basis functions [24]. The mass-weighted normal mode position $\hat{Q}_{\nu}=\sqrt{1/2\omega_{\nu}}(\hat{b}^{\dagger}_{\nu}+\hat{b}_{\nu})$ and the conjugate momentum $\hat{P}_{\nu}=i\sqrt{\omega_{\nu}/2}(\hat{b}^{\dagger}_{\nu}-\hat{b}_{\nu})$ operators for mode $\nu$ can be expressed in terms of the bosonic ladder operators $\{\hat{b}_{\nu}\}$ and $\{\hat{b}^{\dagger}_{\nu}\}$ where $\omega_{\nu}$ is a reference harmonic frequency of normal mode $\nu$. Therefore, the terms depending on the normal mode coordinates in Eq. (3) can be expressed in terms of the bosonic ladder operators by using a Taylor series expansion about a reference geometry $\mathbf{Q}_{0}$ up to a reasonable order.
@@
For fixed Taylor expansion orders, $\omega_{\nu}$ and $\mathbf{Q}_{0}$ only affect the accuracy of our approach by the number of harmonic eigenstates required to describe the wavefunction. In other words, a quantum simulator with lower oscillator noise can tolerate lower-accuracy parameters, including approximations derived from the electronic integral Taylor expansions. For practical cases in the near term, we expect that density functional theory will provide an ideal balance between accuracy and classical computational cost when finding $\mathbf{Q}_{0}$ and $\omega_{\nu}$, and Kohn-Sham orbitals provide a compact spin orbital basis [47].
@@
In general, the orbital vibronic coupling functions $d_{\nu,pq}$ and $g_{\nu,pq}$ can be extremely localized and non-analytical at some nuclear geometries. This requires a high-order expansion in terms of the bosonic ladder operators in Eq. (3), which greatly increases the number of terms in the Hamiltonian. One way to reduce the expansion order would be a unitary transformation of orbitals for which the derivative couplings vanish, which must satisfy $\partial\mathbf{c}/\partial Q_{\nu}+\mathbf{d}_{\nu}\mathbf{c}=0$, where $\mathbf{c}$ is the transformation matrix to the resulting set of “diabatic” orbitals, $\eta_{m}=\sum_{l}c_{lm}\psi_{l}$. This transformation is identical in form to the many-body BO state diabatization condition [14]. However, much like multi-electron states in multi-mode systems [48], a set of strictly diabatic orbitals does not exist for an incomplete basis. Instead, because our goal is simply to achieve a low-order expansion in terms of $\mathbf{Q}$, we can use “quasi-diabatic” orbitals where the above strict diabatization condition is not satisfied but all integral coefficients are smooth functions of $\mathbf{Q}$. These orbitals can likewise borrow techniques from BO state diabatization [49].
@@
Figure 1: Illustration of the pre-BO cMQB analog quantum simulation approach for molecular vibronic dynamics using trapped-ion architectures in comparison with the BO framework. (a) The electronic wavefunction is given in terms of spin orbitals. Each diabatic spatial orbital, which depends on the nuclear positions, yields two spin orbitals with opposite spins: $|\phi_{1}\rangle$ and $|\phi_{3}\rangle$ (green), and $|\phi_{2}\rangle$ and $|\phi_{4}\rangle$ (yellow). Only the motion of a single ion (grey) is shown. (b) Molecular vibronic dynamics are simulated based on the pre-BO representation with interactions among electronic (cyan) and nuclear (grey) degrees of freedom. (c) In this work, we propose a coupled multi-qubit boson approach which maps occupation numbers $\{n_{p}\}$ of spin orbitals $\{|\phi_{p}\rangle\}$ to qubits and nuclear motions to bosonic modes coupled to the qubits, and thus can perform pre-BO dynamics quantum simulations with a complete electronic basis within a given orbital basis set. (d) Existing quantum/classical algorithms rely on BO states $\{|\Phi_{i}\rangle\}$ with potential energy surfaces $\{|E_{i}\rangle\}$ found by solving the electronic time-independent Schrödinger equation using the orbital basis at fixed nuclear positions prior to the simulation. (e) Molecular vibronic dynamics are solved in the BO framework by propagating the nuclear wavefunction (red and blue densities) on a truncated set of coupled BO states. The truncation leads to a deviation from exact results shown in (b).
II.2 Analog quantum simulation